Can someone help me on this problem?
I need to find the constant of variation k
m varies directly with n; m = 144 when n = 8
Thanks,
Ashley

Answer by ankor@dixie-net.com(22740)

:
m varies directly with n; m = 144 when n = 8
:
The equation:
m = n*k
:
Subsitute for m & n, find k
144 = 8k
144/8 = k
k = 18
:
Check it when n = 8
m = 8 * 18
m = 144